AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  AIDA  ELOG logo
Message ID: 617     Entry time: Wed May 31 08:00:26 2017
Author: NN, CB 
Subject: Wednesday 31 May 16:00 - 24:00 
16:30  Check online data ok (R1_533) see attachment 1

16:39  BigRips and BRIKEN stop their runs
       AIDA on file R1_550

16:43  Beam is off

16:47  All degraders completely removed 
       Beam on 
       current AIDA run R1_551

16:48  BRIKEN start run 40
       BigRIPS start run 1041
       AIDA on file R1_551

16:57  BigRIPS and BRIKEN stop their runs
       AIDA on file R1_554
       high energy correlation between x and y attachment 2 

16:57  Beam is off
       Degrader back to its place, thickness 14 mm

17:00  Checked chiller level - 50% 

17:01  Beam is on

17:02  BRIKEN start run 41
       BigRIPS start run 1042
       AIDA on file R1_555

17:16  FEE's temperature normal
       statistics normal
       system wide checks ok 
       leakage current and bias ok see attachment 3
       merger and tapeserver ok

17:17  high energy correlation between x and y attachment 4 
       current AIDA run R1_558

18:02  Beam is off
       current AIDA file Run R1_570
       BigRIPS and BRIKEN stop their runs

       RH = 41.3%, T=24.3 C, DP=10.3 C
       Degrader now 5 (in F7) + 6 (in F11) + 3 (in F11) = 14 mm
       Refilled chiller - >95%

18:42  Beam is on

18:57  The optimization of beam tuning is going on       

19:20  Optimization of beam done

19.28 BRIKEN start run 42
      BigRIPS start run 1053
      AIDA on file R1_588

      BNC PB-4
      Amplitude 90,000
      x5 attenuator IN
      t_r 100ns tau_d 50us
      - polarity
      frequency 25Hz range
      delay MIN, tail pulse, INT ref

      ASIC settings 2017Mar27-16.12.39

      84Zn degrader setting - 5mm (in F7) + 6mm (in F11) + 3mm (in F11) = 14 mm 

19.41 Detector biases; leakage currents OK - see attachments 5 & 6
      FEE64 Temperatures OK - see attachment 7
      good events statistics OK - see attachment 8
      No warnings/errors reported by merger since DAQ GO - see attachment 9

      *all* system wide checks OK *except* ADC calibration of nnaida1, 21 and nnaida1 clock status (status=6)

20.19 DAQ stops 
      AIDA file RIBF127R1/R1_601
      
      analysis of R1_601 dead time <0.5% - see attachment 10
      online analysis - ADC data synchronised - see attachment 12 

      online analysis per FEE64 HEC x versus y - attachment 11 mid-run - all SortSas spectra cleared -
attachment 13 end of run
      why does x-y distribution appear to change?
      distributions centred OK but now broader in x

20.20-20.39 various variable degrader settings to establish approximate degrader setting for 78Ni

20.39 BRIKEN start run 44
      BigRIPS start run 1054
      AIDA on file R1_606

      78Ni degrader setting (to be checked) 5mm Al (fixed F7) + 9mm Al (fixed F11) + 1mm Al (variable F11) = 15mm Al

      Offline sort of R1_602 for rates

 *** DSSSD # 1 count:    156234 old count:    115866 dt:      1.04 s  LEC rate:  38664.03 Hz
 *** DSSSD # 2 count:    205986 old count:    152716 dt:      1.04 s  LEC rate:  51021.42 Hz
 *** DSSSD # 3 count:     72980 old count:     54270 dt:      1.04 s  LEC rate:  17920.23 Hz
 *** DSSSD # 4 count:     63677 old count:     45811 dt:      1.04 s  LEC rate:  17111.86 Hz
 *** DSSSD # 5 count:     24253 old count:     17204 dt:      1.04 s  LEC rate:   6751.45 Hz
 *** DSSSD # 6 count:     48843 old count:     34242 dt:      1.04 s  LEC rate:  13984.68 Hz
 *** DSSSD # 1 count:        75 old count:        52 dt:      1.04 s  HEC rate:     22.03 Hz
 *** DSSSD # 2 count:       114 old count:        82 dt:      1.04 s  HEC rate:     30.65 Hz
 *** DSSSD # 3 count:       118 old count:        78 dt:      1.04 s  HEC rate:     38.31 Hz
 *** DSSSD # 4 count:       107 old count:        67 dt:      1.04 s  HEC rate:     38.31 Hz
 *** DSSSD # 5 count:        54 old count:        30 dt:      1.04 s  HEC rate:     22.99 Hz
 *** DSSSD # 6 count:         4 old count:         4 dt:      1.04 s  HEC rate:      0.00 Hz

20.50 all online histograms zero'd

21:21 analysed run from R1_588 to R1_601, couldn't find any timewraps. 

21:41  BigRips and BRIKEN stop their runs
       AIDA on file R1_620

21:44  BRIKEN start run 45
       BigRIPS start run 1055
       AIDA on file R1_621

22:04  FEE's temperature normal
       statistics normal
       system wide checks ok 
       leakage current and bias ok see attachment 14
       merger and tapeserver ok

22:41  BigRips and BRIKEN stop their runs
       AIDA on file R1_634

22:43  BRIKEN start run 46
       BigRIPS start run 1056
       AIDA on file R1_635

Offline sort of R1_635 for rates

 *** data items:    3524688 ( 108739.47 Hz)
 *** ADC events:     372675 (  11497.32 Hz)
 *** time warps:          0 (      0.00 Hz)
 *** DSSSD # 1 count:    160274 old count:    151136 dt:      3.43 s  LEC rate:  46708.84 Hz
 *** DSSSD # 2 count:    216907 old count:    205441 dt:      3.43 s  LEC rate:  63213.47 Hz
 *** DSSSD # 3 count:     74841 old count:     71214 dt:      3.43 s  LEC rate:  21811.00 Hz
 *** DSSSD # 4 count:     67870 old count:     63212 dt:      3.43 s  LEC rate:  19779.44 Hz
 *** DSSSD # 5 count:     25688 old count:     23773 dt:      3.43 s  LEC rate:   7486.28 Hz
 *** DSSSD # 6 count:     52420 old count:     48302 dt:      3.43 s  LEC rate:  15276.82 Hz
 *** DSSSD # 1 count:        73 old count:        69 dt:      3.43 s  HEC rate:     21.27 Hz
 *** DSSSD # 2 count:       137 old count:       131 dt:      3.43 s  HEC rate:     39.93 Hz
 *** DSSSD # 3 count:       141 old count:       130 dt:      3.43 s  HEC rate:     41.09 Hz
 *** DSSSD # 4 count:       141 old count:       130 dt:      3.43 s  HEC rate:     41.09 Hz
 *** DSSSD # 5 count:        47 old count:        41 dt:      3.43 s  HEC rate:     13.70 Hz
 *** DSSSD # 6 count:         9 old count:         9 dt:      3.43 s  HEC rate:      2.62 Hz

23:20 Check online data ok (R1_642) see attachment 15

23:54  BigRips and BRIKEN stop their runs
       AIDA on file R1_652

*************June 1st********** 

00:03  BRIKEN start run 47
       BigRIPS start run 1058
       AIDA on file R1_654
   
       BigRIPS not sending synchronizing pulse

00:09  BigRIPS and BRIKEN stop their run
Attachment 1: analyser_533.txt  10 kB  Uploaded Wed May 31 09:33:21 2017  | Show | Hide all | Show all
Attachment 2: 310517_energycorrelation1657.png  61 kB  Uploaded Wed May 31 10:00:44 2017  | Hide | Hide all | Show all
310517_energycorrelation1657.png
Attachment 3: 310517_bias1716.png  33 kB  Uploaded Wed May 31 10:21:16 2017  | Hide | Hide all | Show all
310517_bias1716.png
Attachment 4: 310517_energycorrelation1717.png  74 kB  Uploaded Wed May 31 10:22:14 2017  | Hide | Hide all | Show all
310517_energycorrelation1717.png
Attachment 5: 1.png  12 kB  Uploaded Wed May 31 12:43:28 2017  | Hide | Hide all | Show all
1.png
Attachment 6: 2.png  12 kB  Uploaded Wed May 31 12:43:28 2017  | Hide | Hide all | Show all
2.png
Attachment 7: 3.png  134 kB  Uploaded Wed May 31 12:43:28 2017  | Show | Hide all | Show all
Attachment 8: 4.png  92 kB  Uploaded Wed May 31 12:43:28 2017  | Hide | Hide all | Show all
4.png
Attachment 9: 5.png  313 kB  Uploaded Wed May 31 12:43:28 2017  | Hide | Hide all | Show all
5.png
Attachment 10: R1_601.txt  10 kB  Uploaded Wed May 31 13:24:49 2017  | Show | Hide all | Show all
Attachment 11: 6.png  63 kB  Uploaded Wed May 31 13:27:27 2017  | Hide | Hide all | Show all
6.png
Attachment 12: 7.png  78 kB  Uploaded Wed May 31 13:27:34 2017  | Hide | Hide all | Show all
7.png
Attachment 13: 8.png  49 kB  Uploaded Wed May 31 13:29:29 2017  | Hide | Hide all | Show all
8.png
Attachment 14: 310517_bias2204.png  32 kB  Uploaded Wed May 31 15:07:01 2017  | Hide | Hide all | Show all
310517_bias2204.png
Attachment 15: analyser_642.txt  10 kB  Uploaded Wed May 31 16:21:43 2017  | Show | Hide all | Show all
ELOG V3.1.4-unknown